OpenStudy (anonymous):

Rachel is purchasing a condominium and is financing $485,000 with a 30-year 6/1 ARM at 4.95% with a 1/10 cap structure. What will her payments be at the beginning of year 7? $2845.65 $2588.78 $2598.90 $2326.22

OpenStudy (amistre64):

gonna have to restructure it a few times ....

OpenStudy (amistre64):

6 years at the usual fixed rate restructure for the 6th year and the remaining time at +1% restructure again for the 7th year and the remaining time at +2%

OpenStudy (amistre64):

excel is nice for this run .... or you can split it up for the calcs

OpenStudy (amistre64):

Po = 485000k^(12*30)(1-k)/(1-k^(12*30)), k=1+.0495/12 B6 = 485000k^(12*6) - Po (1-k^(12*6))/(1-k), k=1+.0495/12 rework new payments with 30-6 years left at .0595

OpenStudy (amistre64):

P1 = 435809k^(12*24)(1-k)/(1-k^(12*24)), k=1+.0595/12 B7 = 435809k^(12) - Po (1-k^(12))/(1-k), k=1+.0595/12 http://www.wolframalpha.com/input/?i=435809k%5E%2812%29+-+%28435809k%5E%2812*24%29%281-k%29%2F%281-k%5E%2812*24%29%29%29+%281-k%5E%2812%29%29%2F%281-k%29%2C+k%3D1%2B.0595%2F12 so we resturcture for payments using 427364 and .0695 interest for 23 years

OpenStudy (amistre64):

P2 = 427364k^(12*23)(1-k)/(1-k^(12*23)), k=1+.0695/12 http://www.wolframalpha.com/input/?i=427364k%5E%2812*23%29%281-k%29%2F%281-k%5E%2812*23%29%29%2C+k%3D1%2B.0695%2F12 so about 3106 is what i get .. and some errors might have crept in

OpenStudy (anonymous):

is that ur final?

OpenStudy (anonymous):

because i dont have any of those choices

OpenStudy (amistre64):

yeah, but id rather run an excel on it

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

just let me know as soon as u know

OpenStudy (anonymous):

thanks :)

OpenStudy (amistre64):

i might have gone a year to far to begin with 2588 is the payments for the intial term first 6 years, then it bumps by 1 more year, not 2 years so 2845 looks fine

OpenStudy (anonymous):

okay could u help with 3 more questions please :)

OpenStudy (anonymous):

Please please (:

OpenStudy (anonymous):

Brent and Amanda are purchasing a house with a 30-year, 4/1 ARM for $395,000 at 4.65% with a 3/10 cap structure. What will the difference in payments be from year 4 to year 5? $1,282.03 $883.47 $576.77 $686.49

OpenStudy (amistre64):

ive already given you all the tools needed to assess the next one determine the payments for the initial period, and subtract the payment for the next resturcture to find the difference

OpenStudy (anonymous):

A bank advertises a 3/1 ARM at 4.65% with a 3/9 cap. What is the maximum interest rate that can be charged during the fifth year? 3% 4.65% 9% 10.65%